# Styx: Transactional Stateful Functions on Streaming Dataflows

### v2

@@ 7. Fault Tolerance
To ensure atomicity at the level of workflow execution, existing SFaaS systems perform fine-grained fault tolerance (Zhang et al., 2020; Jia and Witchel, 2021a): each function execution is logged and persisted in a workflow log before the next function is called. This requires a round-trip to the logging mechanism for each function call, which adds significant latency to function execution. Styx opts for a coarse-grained fault tolerance mechanism. Instead of logging each function execution, it adopts a variant of existing checkpointing mechanisms used in streaming dataflow systems (Silvestre et al., 2021; Carbone et al., 2017; Chandy and Lamport, 1985): Styx asynchronously snapshots state, enabling very low-latency function execution. We describe this fault tolerance mechanism below.
7.1. Incremental Snapshots & Recovery in Styx
@@
To this end, instead of taking snapshots periodically by propagating markers across the system’s operators, Styx aligns snapshots with the completion of transaction epochs to take a consistent cut of the system’s distributed state – including the state of the latest committed transactions, the offsets of the message broker and, the sequencer counter (). When the snapshot interval triggers, Styx makes a copy of the current state changes to a parallel thread and persists incremental snapshots asynchronously, allowing Styx to continue processing incoming transactions while the snapshot operation is performed in the background. In combination with a replayable source for the incoming transactions (e.g., a Kafka message queue), this approach guarantees the processing of transactions exactly-once, even under failures. In case of a system failure, Styx rolls back to the epoch of the latest complete snapshot, loads the snapshotted state, rolls back the replayable source’s topic partitions (that are aligned with the Styx operator partitions) to the offsets at the time of the snapshot, loads the sequencer counter, and finally, verifies that the cluster is healthy before executing a new transaction epoch.
@@
Incremental Snapshots & Compaction. Each snapshot stores a collection of state changes in the form of delta maps. A delta map is essentially a hash table that tracks the changes in a worker’s state in a given epoch. When a snapshot is taken, only the delta map containing the state changes of the current epoch is snapshotted. To avoid having to track changes across delta maps, Styx performs periodical compactions where the deltas are merged in the background, as shown in Figure 8. In the worst case, Styx will have to merge one delta map to recover a consistent snapshot in case of a failure.
7.2. Exactly-once Output Guarantees
@@
Another common challenge in the fault tolerance of streaming systems is the one of exactly-once output (Elnozahy et al., 2002) in the presence of failures, which is hard to solve for low-latency use cases. For example, in Apache Flink’s (Carbone et al., 2015) exactly-once output configuration, clients can only retrieve responses after those are persisted in a transactional sink. This arrangement is sufficient for streaming analytics but not for low-latency transactional workloads as discussed previously in Section 6.4.
@@
To solve that, Styx persists the output offsets on the output queue, and during recovery Styx reads the last offset of the message broker, compares it with the offset of the message broker persisted in the snapshot to calculate the batch of transactions for processing, retrieves the TIDs attached in the replies of the output queue, and deduplicates the outputs that have already been sent to the output queue, before it continues processing. Note that this deduplication strategy is based on the fact that TIDs have been assigned deterministically.
7.3. Replication and High Availability
Although Styx achieves fault tolerance by storing snapshots in a replicated fault-tolerant blob store (e.g., Minio / S3), it does not yet support other forms of active replication for high availability scenarios. However, Styx in the future can easily inherit the replication mechanisms of deterministic databases. The design of deterministic transaction protocols, such as Calvin (Thomson et al., 2012), feature state replicas that require no explicit synchronization. First, the sequencer replicas need to agree on the order of execution. This is a relatively cheap synchronization step (using Paxos (Lamport, 2001) or Raft (Ongaro and Ousterhout, 2014)), as it does not require the exchange of state, only transaction IDs. After that, the deterministic sequencing algorithm guarantees that the resulting state will be the same across partition/worker replicas by virtue of all replicas executing state updates in the same order.
@@ 8. Evaluation
In this section, we evaluate Styx by answering the following questions:
How does Styx compare to State-of-the-Art serializable transactional SFaaS systems?
How does Styx perform under skewed workload?
How well does Styx scale?
How much does the snapshotting mechanism affect performance?
@@ 8.1. Setup
Systems Under Test. In the evaluation, we only consider SFaaS systems that provide serializable transactional guarantees. At the time of writing, those are: Beldi (Zhang et al., 2020), Boki (Jia and Witchel, 2021a) and T-Statefun (de Heus et al., 2021b).
Beldi/Boki. Both systems use Nightcore (Jia and Witchel, 2021b) as their function runtime, and both use a variant of two-phase commit and store the data in an AWS-managed DynamoDB with enabled autoscaling.
T-Statefun. maintains the state and the coordination of the two-phase commit protocol within an Apache Flink cluster and sends the relevant state to remote stateless function runtimes for execution. For fault tolerance, it relies on a RocksDB state backend that performs incremental snapshots.
Styx. Styx core is implemented in Python 3.12 and uses coroutines to enable asynchronous concurrent execution. Styx uses ZeroMQ111https://zeromq.org/ for messaging between the workers. The system’s ingress point is another message queue, currently Apache Kafka, with important capabilities. It needs to support replayable partitioned fault-tolerant streams that can align with the partitioning of Styx’s operators. Finally, we use Minio222https://min.io//S3 as a remote persistent store for the incremental snapshots. Unless stated otherwise, the snapshot fault-tolerance mechanism is enabled (10-second intervals) across all experiments shown in this section.
@@
YSCB-T (Dey et al., 2014). We use a variant of YCSB-T (Dey et al., 2014) from T-Statefun’s (de Heus et al., 2021a) evaluation. Each transaction consists of two reads and two writes. The concrete scenario is the following: first, we create 10.000 bank accounts (keys) and perform transactions in which a debtor attempts to transfer credit to a creditor. This transfer is subject to a check on whether the debtor has sufficient credit to fulfill the payment. If not, a rollback needs to be performed. The selection of a relatively small number of keys is deliberate: we want to assess the systems’ ability to sustain transactions under high contention. In addition, for the experiment depicted in Figure 10 (skewed distribution), we select the debtor key based on a uniform distribution and the creditor based on a Zipfian distribution, where we can vary the level of contention by modifying the Zipfian coefficient.
TPC-C (Leutenegger and Dias, 1993). A transactional benchmark that targets OLTP systems is TPC-C (Leutenegger and Dias, 1993). To use it in our evaluation, we had to adapt SQL queries into the SFaaS paradigm, splitting the NewOrder transaction into 20-50 function calls (one call for each item in the NewOrder transaction) and the Payment transaction into eight function calls. TPC-C scales in size/partitions by increasing the number of warehouses represented in the benchmark. While a single warehouse represents a skewed workload (all transactions will hit the same warehouse), increasing the number of warehouses decreases the contention allowing for higher throughput and lower latency. Note that the TPC-C experiments do not include Beldi, Boki or T-Statefun: despite our efforts, the existing systems could not handle the complexity of the TPC-C workload (complex call graph) in order to produce meaningful experimental results for the needs of this paper.
| 0.0 | 0.1 | 0.2 | 0.3 | 0.4 | 0.5 | 0.6 | 0.7 | 0.8 | 0.9 | 0.99 | 0.999 | ||
| Beldi | Abort % | 47.93 | 46.58 | 45.54 | 45.68 | 44.31 | 48.68 | 47.28 | 48.70 | 52.40 | 56.06 | 61.62 | 60.70 |
| CMT TPS | 104 | 108 | 108 | 111 | 111 | 105 | 105 | 95 | 95 | 76 | 76 | 78 | |
| Boki | Abort % | 48.77 | 48.83 | 48.23 | 47.38 | 49.54 | 48.29 | 51.82 | 55.69 | 61.29 | 68.50 | 74.47 | 70.71 |
| CMT TPS | 359 | 358 | 362 | 368 | 353 | 362 | 337 | 310 | 271 | 220 | 179 | 205 |
Resources. For Beldi/Boki, T-Statefun and Styx, we assigned a total of 112 CPUs with 2GBs of RAM per CPU, matching what is presented in the original Boki paper (Jia and Witchel, 2021a). Unless stated otherwise, Styx and T-Statefun are configured to perform incremental snapshots every 10 seconds. All external systems i.e., DynamoDB (Beldi, Boki), Minio, and Kafka (transaction ingress for Styx, T-Statefun) are configured with three replicas for fault tolerance.
External Systems. Boki and Beldi make use of a fully managed and autoscaled DynamoDB instance (configured to use 3 replicas) at AWS. The resources occupied by DynamoDB are additional to the 112 CPUs assigned to Boki and Beldi. Similarly, the resources assigned to Styx and T-Statefun do not include their snapshot store, Minio/S3.
Metrics. We are mainly interested in two important metrics.
Input throughput. represents the number of transactions that are submitted, per second, to the system under test. As the input throughput increases, throughout an experiment, we expect the latency of individual transactions to increase, until aborts start to manifest due to contention or high load.
Latency. represents the time interval between submitting a transaction, and the reported time when the transaction committed/aborted. In Styx and T-Statefun the latency timer starts when a transaction is submitted in the input queue (Kafka). The end of the transaction is when the system reports the transaction as committed/aborted in the output queue. Similarly, in Beldi and Boki the latency is the time since the input gateway has received a transaction, and the time that the gateway reports that the transaction has been committed/aborted.
@@
We first study the latency-throughput tradeoff of all systems. We retain the resources given to the systems constant (112 CPUs) while we progressively increase the input throughput. We monitor the transaction latency. As depicted in Figure 9, Styx outperforms its competitors by at least an order of magnitude. Specifically, Styx achieves a performance improvement of ~20x against T-Statefun, which ranks second, in YCSB-T (Figure 8(a)) in terms of throughput, ~30x against Boki in Deathstar’s travel reservation workload, (Figure 8(b)) and ~35x against Boki in Deathstar’s movie review Figure 8(c)) workload. Finally, in the TPC-C benchmark (Figure 8(d)), which is the most demanding of all, as we increase the input throughput for different number of warehouses, we observe that Styx can handle up to 3K TPS with sub-second 99 percentile latency (100 warehouses).
@@
Styx and T-Statefun executed without any concurrency-related aborts. On the contrary, Beldi and Boki utilize a no-wait-die concurrency control mechanism, which leads to a significant amount of aborts as the throughput increases, making the latency measurements incomparable. On this ground, in Figure 10 we plot the results of Styx and T-Statefun and present the performance of Beldi and Boki including their abort rates in a separate table.
@@
We observe the following: at the highest level of contention ( at ) Styx achieves 2000 TPS, outperforming the rest by ~5-10x in terms of effective throughput (see table in Figure 10), both Beldi and Boki abort more transactions as the level of contention increases (~40-70%), which significantly impacts their effectiveness, and Styx manifests an increase in latency only in high levels of contention () while executing at ~4x higher throughput than the rest.
@@
Systems’ Performance Breakdown. In Figure 11, we show where the systems under test spend their processing time. We use YCSB-T for this purpose since it is the benchmark with the lowest function execution complexity, making it suitable for testing how networking and state access affect performance. We measured the median latency while all the systems were running at 100 TPS for 60 seconds, and averaged the proportions of function execution, networking, and state access across all committed transactions. The key observations are: Styx’s co-location of processing and state led to near-zero state access latency and; Styx’s asynchronous networking allows for very low network latency.
@@
Experiments Takeaway. The rather large performance gains of Styx across all experiments are enabled by the following four properties and design choices: the co-location of processing and state with efficient networking as shown in Figure 11 contrary to the other systems that have to transfer the state to their function execution engines; the asynchronous snapshots with delta maps for fault tolerance compared to the replication of Beldi/Boki and the LSM-tree-based incremental snapshots of T-Statefun; the efficient transaction execution protocol employed in Styx compared to the two-phase commit used by Beldi, Boki, and T-Statefun; and the underlying performance of Styx’s streaming dataflow engine.

### v4

@@ 7. Fault Tolerance
Styx implements a coarse-grained fault tolerance mechanism. Instead of logging each function execution, it adopts a variant of existing checkpointing mechanisms used in streaming dataflow systems (Silvestre et al., 2021; Carbone et al., 2017; Chandy and Lamport, 1985). Styx asynchronously snapshots state and stores it in a replicated fault-tolerant blob store (e.g., Minio / S3), enabling low-latency function execution. We describe Styx’s fault tolerance mechanism below.
7.1. Incremental Snapshots & Recovery
@@
Snapshotting. To this end, instead of taking snapshots periodically by propagating markers across the system’s operators, Styx aligns snapshots with the completion of transaction epochs to take a consistent cut of the system’s distributed state, including the state of the latest committed transactions, the offsets of the message broker, and the sequencer counters (). The minimal information included in the snapshot is where is the number of updates affecting the delta map, and is the fixed number of integers stored for the Kafka offsets and the sequencer variables.
When the snapshot interval triggers, Styx makes a copy of the current state changes to a parallel thread and persists incremental snapshots asynchronously, allowing Styx to continue processing incoming transactions while the snapshot operation is performed in the background. The snapshotting procedure is described in Algorithm 1.
Recovery. In case of a system failure, Styx rolls back to the epoch of the latest completed snapshot, loads the snapshotted state, rolls back the replayable source’s topic partitions (that are aligned with the Styx operator partitions) to the offsets at the time of the snapshot, loads the sequencer counters, and finally, verifies that the cluster is healthy before executing a new epoch. The recovery procedure is described in Algorithm 2.
@@
Incremental Snapshots & Compaction. Each snapshot stores a collection of state changes in the form of delta maps. A delta map is a hash table that tracks the changes in a worker’s state in a given snapshot interval. When a snapshot is taken, only the delta map containing the state changes of the current interval is snapshotted. To avoid tracking changes across delta maps, Styx periodically performs compactions where the deltas are merged in the background, as shown in Figure 8. The cost of compacting is equivalent to the cost of merging two hashmaps with the same key-spaces . The total cost will be , with denoting the number of deltamaps we need to compact.
7.2. Sequencer Recovery
To guarantee determinism, upon recovery, Styx ’s sequencer needs to generate identical sequences as the ones generated between the latest snapshot and failure. The recovery protocol of the sequencer operates as follows: At first, during the snapshot, we store the local counter of each sequencer partition () with its id () and the epoch counter. Additionally, at the start of each epoch, Styx logs the number of transactions contained in that epoch, denoted as epoch size. Logging the epoch sizes is needed due to Styx ’s varying epoch sizes and the sequencer rebalancing scheme (section 5.1). After failure, the recovered sequencer partitions are initialized with the snapshot’s and . Afterward, each partition retrieves from its log all the sizes of all epochs executed since the last snapshot. Finally, after recovery, the sequencer matches the epoch sizes to the ones recorded by the log, leading to the same global sequence observed before failure.
7.3. Exactly-Once Processing
At first, the durable input queue, which acts as a replayable source, allows Styx to replay requests after failures. By rolling back the queue partitions (aligned with system operator partitions) to the respective offsets as recorded in the latest snapshot, Styx can reprocess only the transactions whose state changes are not reflected yet in the snapshot. Transactions committed and early-commit replies stored in the egress can be deduplicated (Section 7.4).
Styx runs each transaction to its completion in a single epoch. A given transaction can execute a large call-graph of functions that can affect the state. If a failure takes place, a transaction’s state effects are restored to the latest snapshot, and the complete transaction is re-executed. As a result, no special attention is required to ensure that remote function calls are executed exactly-once, except for resetting all TCP channels between Styx’s workers after recovery.
Lemma 7.1.
The state mutations of committed transactions in Styx are reflected exactly-once, even upon failure.
Proof.
Let denote the state of the system at time . denotes the durable input queue at time that holds all requests to be processed. We assume that the input queue operates as FIFO and requests are deterministic. Each will be sequenced as a transaction by a deterministic sequencer, where are the state updates and are the function calls of the transaction. We assume that happens atomically and are also reflected once, given the use of a reliable communication protocol. Given the same initial state and input from , it always produces the same state transition ,which means . The execution of a transaction is deterministic.
At any time , the state of the system reflects all transactions in that have been fully executed and committed. Accordingly, the state ignores partially executed or in-progress transactions in . We denote the latest durable snapshot taken up to time , as where corresponds to the offsets of the first request , and last request of the input queue to be processed up to time . Upon failure, a subset of , will contain successfully committed transactions and a subset will contain aborted transactions such that . In order to recover from a failure, is rolled back to from as we persist the offsets of our input queue. Transactions in are replayed in the original order from offset to offset of our input queue. This is ensured by the FIFO queue and the deterministic sequencer. After processing the input transactions, includes requests already reflected in , and includes pending requests. Since reflects and , the replay and processing ensure: . Thus, the effects of all transactions will be reflected in the state exactly-once, even after failure. ∎
7.4. Exactly-Once Output
@@
A common challenge in the fault tolerance of streaming systems is that of the exactly once output (Elnozahy et al., 2002; Fragkoulis et al., 2024) in the presence of failures, which is hard to solve for low-latency use cases. For example, in Apache Flink’s (Carbone et al., 2015) exactly-once output configuration, clients can only retrieve responses after those are persisted in a snapshot or a transactional sink. This arrangement is sufficient for streaming analytics but not for low-latency transactional workloads, as discussed previously in Section 6.4.
@@
To solve that, during recovery, Styx: reads the last offset of the egress topic, compares it with the output offset persisted in the snapshot, determining for which transactions the clients have already received replies, retrieves the TIDs attached in those replies, and does not send a reply again to the egress topic for those transactions. Note that this deduplication strategy is based on the fact that TIDs have been assigned deterministically.
7.5. Addressing Non-Deterministic Functions
As discussed in Section 7.1 Styx’s recovery mechanism is based on deterministic replay. To this end, Styx requires that the functions authored by developers are also deterministic, i.e., replaying the same function multiple times, using the same inputs and database state, should yield the same results. However, one can achieve determinism even in the presence of non-deterministic logic inside functions, such as randomness (e.g., random numbers/sampling) or calls to external systems (e.g., calling an external database or API). Styx can follow the approach of existing systems (e.g., Temporal (Temporal, 2025), Clonos (Silvestre et al., 2021)). In the following, we explain how this can be achieved.
Randomness. To retain determinism in the case of randomness, Styx can use an external fault tolerant write-ahead log (WAL) to log the random number along with the TID. Thus, in the case of failure and replay, Styx can use the logged random number, essentially making the function call deterministic during replay.
Calls to External Systems. As illustrated in Figure 9, an interaction with an external system needs to consider three critical points to maintain determinism. Styx assumes that the external system supports idempotency (IETF, 2025), meaning that if a call is made twice with the same idempotency key, the effects on the external system’s state and its return value will remain the same. In Styx needs to log the idempotency key and the TID in the WAL before calling the external system. If the external system produces a response ( ), Styx can store it in the WAL and retrieve it from there in case of replay. Finally, when Styx completes a snapshot ( ), it can also clear the WAL for garbage collection since the prior entries are not needed.
Finally, Styx could mask those operations behind an API that exposes the following functionality such as system_x.random for random number generation and system_x.call_external for external system calls.
@@ 8. Evaluation
We evaluate Styx by answering the following questions:
(Section 8.2) How does Styx compare to State-of-the-Art serializable transactional SFaaS systems?
(Section 8.2) How does Styx perform under skewed workload?
(Section 8.3) How well does Styx scale?
(Section 8.4) Does the snapshotting mechanism affect performance?
@@ 8.1. Setup
Systems Under Test. In the evaluation, we include SFaaS systems that provide serializable transactional guarantees. Those are:
Beldi (Zhang et al., 2020)/Boki (Jia and Witchel, 2021a). Both systems use a variant of two-phase commit and Nightcore (Jia and Witchel, 2021b) as their function runtime and store their data in DynamoDB. Additionally, Boki is deployed with the latest improvements of Halfmoon (Qi et al., 2023).
T-Statefun (de Heus et al., 2021). T-Statefun maintains the state and the coordination of the two-phase commit protocol within an Apache Flink cluster and ships the relevant state to remote stateless functions for execution. For fault tolerance, it relies on a RocksDB state backend that performs incremental snapshots.
Styx. Styx is implemented in Python 3.12 and uses coroutines to enable asynchronous concurrent execution. Apache Kafka is used as an ingress/egress and Minio/S3 (Min, [n. d.]) as a remote persistent store for Styx’s incremental snapshots. Finally, Styx is a standalone containerized system that works on top of Docker and Kubernetes for ease of deployment.
YSCB-T (Dey et al., 2014). We use a variant of YCSB-T (Dey et al., 2014) where each transaction consists of two reads and two writes. The concrete scenario is as follows: First, we create 10.000 bank accounts (keys) and perform transactions in which a debtor attempts to transfer credit to a creditor. This transfer is subject to a check on whether the debtor has sufficient credit to fulfill the payment. If not, a rollback needs to be performed. The selection of a relatively small number of keys is deliberate: we want to assess the systems’ ability to sustain transactions under high contention. In addition, for the experiment depicted in Figure 11 (skewed distribution), we select the debtor key based on a uniform distribution and the creditor based on a Zipfian distribution, where we can vary the level of contention by modifying the Zipfian coefficient.
TPC-C (Leutenegger and Dias, 1993). The prime transactional benchmark targeting OLTP systems is TPC-C (Leutenegger and Dias, 1993). In our evaluation, we employ the NewOrder and Payment transactions, and we had to rewrite them into the SFaaS paradigm, splitting the NewOrder transaction into 20-50 function calls (one call for each item in the NewOrder transaction) and the Payment transaction into 8 function calls. TPC-C scales in size/partitions by increasing the number of warehouses represented in the benchmark. While a single warehouse represents a skewed workload (all transactions will hit the same warehouse), increasing the number of warehouses decreases the contention, allowing for higher throughput and lower latency. Note that the TPC-C experiments do not include Beldi, Boki, or T-Statefun because they do not support it.
@@
Resources. For Beldi/Boki, T-Statefun and Styx, we assigned a total of 112 CPUs with 2GBs of RAM per CPU, matching what is presented in the original Boki paper (Jia and Witchel, 2021a). Additionally, throughout all the evaluation scenarios, the data fit in memory across all systems. Unless stated otherwise, Styx and T-Statefun are configured to perform incremental snapshots every 10 seconds. All external systems, i.e., DynamoDB (Beldi, Boki), Minio, and Kafka (Styx, T-Statefun), are configured with three replicas for fault tolerance.
External Systems. Boki and Beldi use a fully managed DynamoDB instance at AWS, which does not state the amount of resources it occupies and is additional to the 112 CPUs assigned to Boki and Beldi. Similarly, the resources assigned to Minio/S3 (Styx and T-Statefun) are not accounted for.
Metrics. Our goal is to observe systems’ behavior, measured by their latency while varying the input throughput.
Input throughput represents the number of transactions submitted per second to the system under test. As the input throughput increases during an experiment, we expect the latency of individual transactions to increase until aborts start to manifest due to contention or high load.
Latency represents the time interval between submitting a transaction and the reported time when the transaction is committed/aborted. In Styx and T-Statefun, the latency timer starts when a transaction is submitted in the input queue (Kafka) and stops when the system reports the transaction as committed/aborted in the output queue. Similarly, in Beldi and Boki, the latency is the time since the input gateway has received a transaction and the time that the gateway reports that the transaction has been committed/aborted.
@@
We first study the latency-throughput tradeoff of all systems. We retain the resources given to the systems constant (112 CPUs) while progressively increasing the input throughput. We measure the transaction latency. As depicted in Figure 10, Styx outperforms its baseline systems by at least an order of magnitude. Specifically, in YCSB-T (Figure 10(a)), Styx achieves a performance improvement of ~20x in terms of throughput against T-Statefun, which ranks second. In addition, Styx outperforms Boki by ~30x in Deathstar’s travel reservation workload (Figure 10(b)) and by ~35x in Deathstar’s movie review Figure 10(c)) workload. Finally, in the TPC-C benchmark (Figure 10(d)), which requires a large number of function calls per transaction (20-50), we observe that Styx’s performance improves as we increase the input throughput for different numbers of warehouses, reaching up to 3K TPS with sub-second 99 percentile latency (100 warehouses).
@@
Aborts & Throughput. Beldi and Boki follow a no-wait-die concurrency control approach, which leads to a significant amount of aborts as the throughput increases. Styx and T-Statefun do not use such a transaction abort mechanism. Instead, they execute all transactions to completion. This difference in handling transactions under high load makes the latencies across systems hard to compare. For this reason, in Figure 11, we plot the results of Styx and T-Statefun and present the performance of Beldi and Boki in a separate table (Table 2), alongside their abort rates.
0.0 | 0.2 | 0.4 | 0.6 | 0.8 | 0.9 | 0.99 | 0.999 | ||
Beldi | Abort % | 47.93 | 45.54 | 44.31 | 47.28 | 52.40 | 56.06 | 61.62 | 60.70 |
CMT TPS | 104 | 108 | 111 | 105 | 95 | 76 | 76 | 78 | |
Boki | Abort % | 48.77 | 48.23 | 49.54 | 51.82 | 61.29 | 68.50 | 74.47 | 70.71 |
CMT TPS | 359 | 362 | 353 | 337 | 271 | 220 | 179 | 205 |
@@
We observe the following: at the highest level of contention ( at ) Styx achieves at least 2000 TPS, outperforming the rest by ~5-10x in terms of effective throughput, both Beldi and Boki (that run at their maximum sustainable throughput) abort more transactions as the level of contention increases (~40-70%), which significantly impacts their effectiveness as shown in Table 2, and Styx shows an increase in latency only in high levels of contention () while executing at ~4x higher throughput than the rest.
@@
Runtime Breakdown. In Table 3, we show where the systems under test spend their processing time. We use YCSB-T for this purpose since it is the only benchmark supported by all the systems (Section 8.1). We measured the median latency while all the systems were running at 100 TPS for 60 seconds and averaged the proportions of function execution, networking, and state access across all committed transactions. The key observations are: Styx’s co-location of processing and state led to minimal state access latency, and Styx’s asynchronous networking allows for lower network latency.
System | Function Execution | Networking | State Access |
Styx | 0.34ms - 2.2% | 14.33ms - 95.6% | 0.32ms - 2.2% |
Boki | 1.1ms - 3.3% | 16.1ms - 49% | 15.68ms - 47.7% |
T-Statefun | 2.76ms - 2.2% | 92.12ms - 74.3% | 29.11ms - 23.5% |
Beldi | 1.01ms - 0.7% | 56.58ms - 38.4% | 89.57ms - 60.9% |
@@
Takeaway. The rather large performance advantages of Styx across all experiments are enabled by the following three properties and design choices: the co-location of processing and state with efficient networking as shown in Table 3, contrary to the other systems that have to transfer the state to their function execution engines; the asynchronous snapshots with delta maps for fault tolerance compared to the replication of Beldi/Boki and the LSM-tree-based incremental snapshots of T-Statefun; the efficient transaction execution protocol employed in Styx compared to the two-phase commit used by Styx’s competition.